OpenStudy (anonymous):

which of the following is a factor of x^3+x^2-6x? A. x-1 B. x-6 C. x+3 D.X+2 i know its C but i dnt understand it

OpenStudy (anonymous):

First factor out an x as a gcf, what do you get?

OpenStudy (anonymous):

After, factor the remaining equation

OpenStudy (amistre64):

if its a factor, then they will bot have the same zeros ....

OpenStudy (anonymous):

yeah so 3 goes both in 6 and 3

OpenStudy (anonymous):

what do you mean by zeros?

OpenStudy (anonymous):

I dont believe shes there yet.

OpenStudy (amistre64):

\[\frac{(x-h)(x-k)}{(x-h)}=\frac{0}{0},~when~x=h \]

OpenStudy (anonymous):

im confused